Lori Cook has developed the following forecasting model:

$$ \hat{y}=45.0+4.50 \mathrm{x} $$

where \(\hat{y}=\) demand for Kool Air conditioners and \(x=\) the outside temperature (degrees Fahrenheit)

a) When the temperature outside is \(70^{\circ} \mathrm{F}\), demand forecast = air conditioners (enter your response as an integer).

b) When the temperature outside is \(80^{\circ} \mathrm{F}\), demand forecas= air conditioners (enter your response as an integer).

c) When the temperature outside is \(90^{\circ} \mathrm{F}\), demand forecast = air conditioners (enter your response as an integer).
Answer #1

y = 45.0 + 4.50 X

a)

Demand forecast when X = 70, Y = 45.0 + 4.50 x 70 = 360

b)

Demand forecast when X = 80, Y = 45.0 + 4.50 x 80 = 405

c)

Demand forecast when X = 90, Y = 45.0 + 4.50 x 90 = 450